# Solar System Constraints on Light Propagation from Higher Derivative Corrections to General Relativity and Implications for Quantum Gravity

### v2

@@ Abstract
While the two derivative action of gravitation is specified uniquely, higher derivative operators are also allowed with coefficients that are not specified uniquely by effective field theory. We focus on a four derivative operator in which the Riemann tensor couples directly to the electromagnetic field $a\,R_{\mu\nu\alpha\beta}F^{\mu\nu}F^{\alpha\beta}$. We compute the corresponding corrections to the Shapiro time delay in the solar system and compare this to data from the Cassini probe. We place an observational upper bound on the coefficient $a$ at 95% confidence $|a|<26\,(1000\,\mbox{km})^{2}$. By way of motivation, we also compare this to a weak gravity conjecture (WGC) prediction of a bound on the coefficients $a,\,b$ of four derivative operators involving the graviton and the photon; this includes the above term $a\,R_{\mu\nu\alpha\beta}F^{\mu\nu}F^{\alpha\beta}$ as well as $b\,F^{4}$. We show that by using the observed value of the $b$ coefficient from measurements of light by light scattering, which arises in the Standard Model from integrating out the electron, the WGC predicted bound for $a$ is $a\lesssim 7.8\,(1000\,\mbox{km})^{2}$. This is consistent with the above observational bound, but is intriguingly close and can be further probed in other observations.
@@ I Introduction
Normally, these inequalities are thought to be robust but very difficult to test experimentally. The most well known inequality from the weak gravity conjecture (WGC) Arkani-Hamed:2006emk is $m\leq\sqrt{2}\,e\,M_{\mbox{\tiny{Pl}}}$, where $m$ is the mass of the lightest charged state, $e$ is the electric coupling, and $M_{\mbox{\tiny{Pl}}}=1/\sqrt{8\pi G_{N}}$ is the Planck mass. If we apply this to the electron with $m_{e}\approx 0.5\,$MeV, electric coupling $e\approx 0.3$ and Planck mass $M_{\mbox{\tiny{Pl}}}\approx 2.4\times 10^{18}\,$eV, this inequality is satisfied by many orders of magnitude. By contrast, the weak gravity conjecture inequality that we will focus on in this work is surprisingly testable, as we will explain. This will come with the very large caveat that to be near the inequality, the cut off on the effective theory will be very low, requiring new physics to enter at very low scales. While such low scale UV cut offs do not sound (to us) very plausible, we think it is still worth subjecting any claims from quantum gravity to the test, regardless of how weak the tests may be.
@@ III Measurements and
In the approximate vacuum of the solar system away from the sun, we have $R_{\mu\nu}\approx 0$; so the $a^{\prime}$ term will not play a significant role. While we still have a non-zero $R_{\mu\nu\alpha\beta}$ from the sun; so the $a$ term will be essential. The above WGC inequality only refers to the sum, but barring some highly unexpected cancellation [^1] For example, integrating out the electron at 1-loop leads to $\Delta a^{\prime}=-13\Delta a$ Drummond:1979pp . Also, the coefficients should obey $a^{\prime}=-4a$ to avoid higher derivative equations of motion. Or we should have $a^{\prime}=-2a$ to form the Weyl tensor. So in all known examples to us, there is no cancellation of the form $a=-a^{\prime}$., the WGC bound can be interpreted roughly as $a\lesssim 7.8(1000\,\mbox{km})^{2}$. So we will be effectively referring to just $a$ from now on. In fact one can go further (see Refs. Hamada:2018dde ; Alberte:2020bdz ): by performing some field redefinitions, one can take all the $a,\,a^{\prime},\,a^{\prime\prime}$ terms to be combined into the form $-a\,W_{\mu\nu\alpha\beta}F^{\mu\nu}F^{\alpha\beta}$, where $W_{\mu\nu\alpha\beta}$ is the Weyl tensor (giving $a^{\prime}=-2a$ and $a^{\prime\prime}=a/3$); this makes this a more direct photon-graviton coupling. (Terms of the form $R_{\mu\nu}F^{\mu\alpha}F^{\nu}_{\,\,\alpha}$ and $RF^{2}$ are not true graviton-photon couplings in a sense: This is because the Einstein equations tell us to leading order $R_{\mu\nu}$ is determined by $T_{\mu\nu}$, which in turn is determined by a bilinear in $F_{\mu\nu}$ the electromagnetic field strength in vacuum. Upon substitution this means the $a^{\prime},\,a^{\prime\prime}$ terms, to leading order, play a redundant role to the terms quartic in $F_{\mu\nu}$, which we are already including.) Alternatively, one can set $a^{\prime}=-4a$ and $a^{\prime\prime}=a$ to achieve two derivative equations of motion Horndeski:1976gi ; Barrow:2012ay ; BeltranJimenez:2013btb . In either way, the values of $a^{\prime},\,a^{\prime\prime}$ terms would be uniquely determined by $a$.
@@ IV Altered Light Travel Time
Figure 1: Arrangement in solar system: Dashed line is light path of radio waves between earth and Cassini, $b$ is impact parameter, $r_{1}$ and $r_{2}$ are the distances from sun to earth and sun to the Cassini probe, respectively.
@@
So in order to avoid superluminality, we need $16|a|/b^{2}<$ 1. As described earlier, we will be interested in values of $a$ in the vicinity of $a\sim 10\,(1000\,\mbox{km})^{2}$, while the smallest impact parameter we will study is $b_{\mbox{\tiny{min}}}=1.6\,R_{\odot}\approx 1.1\times 10^{6}\,$km. Hence we have $6|a|/b_{\mbox{\tiny{min}}}^{2}\sim 5\times 10^{-5}$, ensuring that the behavior is always subluminal (away from the maximum speed correction, the corrections from this new operator are even smaller). We note that other discussions of superluminality in related contexts, includes Refs. Daniels:1993yi ; Dittrich:1998fy ; Novello:1999pg ; Shore:2007um ; Hollowood:2008kq ; Hollowood:2015elj ; Goon:2016une ; Hertzberg:2017abn ; deRham:2020zyh .
@@ V Cassini Probe Data
So we see that the higher derivative correction is a fractional correction of $-8\varepsilon a/b^{2}$.
@@ VI Likelihood Results
In the figure, we have also indicated the bound given by the WGC of $a\lesssim 7.8(1000\,\mbox{km})^{2}$ as the vertical dashed green lines. Assuming $a+a^{\prime}$ is positive, then we should remain within this range $|\bar{a}|\lesssim 7.8(1000\,\mbox{km})^{2}$ or one of the polarizations would be inconsistent. (If $a+a^{\prime}$ is negative, we have no direct bound from WGC).
@@ VII Discussion
We have found an observational bound on the $R_{\mu\nu\alpha\beta}F^{\mu\nu}F^{\alpha\beta}$ coefficient of $|a|<26\,(1000\,\mbox{km})^{2}$. An important point is that the light propagation is polarization dependent. Ideally one would separate out the two polarizations and analyze each individually. However the data is not currently presented in this form. Instead we have exploited the fact that each individual data point has not shown a splitting into pairs (one raised and one lowered), or an enhanced variance, as would be expected if the higher dimension operator is altering the propagation, and hence we can still infer a bound on $a$. Our statistical analysis is a first pass, given that we do not have explicit polarization data. It would be useful in future work to search explicitly for a possible deviation in the behavior of each polarization. The two relevant polarizations are linear polarizations; one polarized orthogonal to the solar system and the other polarized parallel to the solar system. This may lead to a significantly improved bound (or a possible detection).

### v3

@@ Abstract
While the two derivative action of gravitation is specified uniquely, higher derivative operators are also allowed with coefficients that are not specified uniquely by effective field theory. We focus on a four derivative operator in which the Riemann tensor couples directly to the electromagnetic field $a\,R_{\mu\nu\alpha\beta}F^{\mu\nu}F^{\alpha\beta}$. We compute the corresponding corrections to the Shapiro time delay in the solar system and compare this to data from the Cassini probe. We place an observational upper bound on the coefficient $a$ at 95% confidence $|a|<26\,(1000\,\mbox{km})^{2}$. By way of motivation, we also compare this to a weak gravity conjecture (WGC) prediction of a bound on the coefficients a,ba,\,b of four derivative operators involving the graviton and the photon; this includes the above term $a\,R_{\mu\nu\alpha\beta}F^{\mu\nu}F^{\alpha\beta}$ as well as $b\,F^{4}$. We show that by using the observed value of the $b$ coefficient from measurements of light by light scattering, which arises in the Standard Model from integrating out the electron, the WGC predicted bound for $a$ is $a\lesssim 7.8\,(1000\,\mbox{km})^{2}$. This is consistent with the above observational bound, but is intriguingly close and can be further probed in other observations.
@@ I Introduction
Normally, these inequalities are thought to be robust but very difficult to test experimentally. The most well known inequality from the weak gravity conjecture (WGC) Arkani-Hamed:2006emk is $m\leq\sqrt{2}\,e\,M_{\mbox{\tiny{Pl}}}$, where $m$ is the mass of the lightest charged state, $e$ is the electric coupling, and $M_{\mbox{\tiny{Pl}}}=1/\sqrt{8\pi G_{N}}$ is the Planck mass. If we apply this to the electron with $m_{e}\approx 0.5\,$MeV, electric coupling $e\approx 0.3$ and Planck mass $M_{\mbox{\tiny{Pl}}}\approx 2.4\times 10^{18}\,$GeV, this inequality is satisfied by many orders of magnitude. By contrast, the weak gravity conjecture inequality that we will focus on in this work is surprisingly testable, as we will explain. This will come with the very large caveat that to be near the inequality, the cut off on the effective theory will be very low, requiring new physics to enter at very low scales. While such low scale UV cut offs do not sound (to us) very plausible, we think it is still worth subjecting any claims from quantum gravity to the test, regardless of how weak the tests may be.
@@ III Measurements and
In the approximate vacuum of the solar system away from the sun, we have $R_{\mu\nu}\approx 0$; so the $a^{\prime}$ term will not play a significant role. While we still have a non-zero $R_{\mu\nu\alpha\beta}$ from the sun; so the $a$ term will be essential. The above WGC inequality only refers to the sum, but barring some highly unexpected cancellation [^1] For example, integrating out the electron at 1-loop leads to $\Delta a^{\prime}=-13\Delta a$ Drummond:1979pp . Also, the coefficients should obey $a^{\prime}=-4a$ to avoid higher derivative equations of motion. Or we should have $a^{\prime}=-2a$ to form the Weyl tensor. So in all known examples to us, there is no cancellation of the form $a=-a^{\prime}$., the WGC bound can be interpreted roughly as $a\lesssim 7.8(1000\,\mbox{km})^{2}$. So we will be effectively referring to just $a$ from now on. In fact one can go further (see Refs. Hamada:2018dde ; Alberte:2020bdz ): by performing some field redefinitions, one can take all the $a,\,a^{\prime},\,a^{\prime\prime}$ terms to be combined into the form $a\,W_{\mu\nu\alpha\beta}F^{\mu\nu}F^{\alpha\beta}$, where $W_{\mu\nu\alpha\beta}$ is the Weyl tensor (giving $a^{\prime}=-2a$ and $a^{\prime\prime}=a/3$); this makes this a more direct photon-graviton coupling. (Terms of the form $R_{\mu\nu}F^{\mu\alpha}F^{\nu}_{\,\,\alpha}$ and $RF^{2}$ are not true graviton-photon couplings in a sense: This is because the Einstein equations tell us to leading order $R_{\mu\nu}$ is determined by $T_{\mu\nu}$, which in turn is determined by a bilinear in $F_{\mu\nu}$ the electromagnetic field strength in vacuum. Upon substitution this means the $a^{\prime},\,a^{\prime\prime}$ terms, to leading order, play a redundant role to the terms quartic in $F_{\mu\nu}$, which we are already including.) Alternatively, one can set $a^{\prime}=-4a$ and $a^{\prime\prime}=a$ to achieve two derivative equations of motion Horndeski:1976gi ; Barrow:2012ay ; BeltranJimenez:2013btb . In either way, the values of $a^{\prime},\,a^{\prime\prime}$ terms would be uniquely determined by $a$.
@@ IV Altered Light Travel Time
Figure 1: Arrangement in solar system: Dashed line is light path of radio waves between earth and Cassini, $b$ is impact parameter, $r_{1}$ and $r_{2}$ are the distances from sun to earth and sun to Cassini probe, respectively.
@@
So in order to avoid superluminality, we need $6|a|/b^{2}<1$. As described earlier, we will be interested in values of $a$ in the vicinity of $a\sim 10\,(1000\,\mbox{km})^{2}$, while the smallest impact parameter we will study is $b_{\mbox{\tiny{min}}}=1.6\,R_{\odot}\approx 1.1\times 10^{6}\,$km. Hence we have $6|a|/b_{\mbox{\tiny{min}}}^{2}\sim 5\times 10^{-5}$, ensuring that the behavior is always subluminal (away from the maximum speed correction, the corrections from this new operator are even smaller). We note that other discussions of superluminality in related contexts, includes Refs. Daniels:1993yi ; Dittrich:1998fy ; Novello:1999pg ; Shore:2007um ; Hollowood:2008kq ; Hollowood:2015elj ; Goon:2016une ; Hertzberg:2017abn ; deRham:2020zyh .
@@ V Cassini Probe Data
So we see that the higher derivative correction is a fractional correction of $8\,\varepsilon\,a/b^{2}$.
@@ VI Likelihood Results
Recall that the relation between $a$ and $\bar{a}$ is that $\bar{a}=\varepsilon a$. And the value of $\varepsilon$ is $\varepsilon=+1$ for polarization perpendicular to the plane of Cassini-earth sun, and $\varepsilon=-1$ for polarization parallel parallel to the plane of Cassini-earth-sun. These are orthogonal linear polarizations which are normal modes of this system. One can indeed have $\varepsilon$ between -1 and 1 too, but not for the normal modes. So such other modes are superpositions of the normal modes, and the other values will have non-trivial time dependence. In Fig. (3), the constraints (blue curve) pertain to the variable $\bar{a}$ as the correction to the Shapiro delay is proportional to this. The vertical green lines indicate the WGC bounds, which are proportional to $a$. But since $|\varepsilon|\leq 1$, we can say that the WGC bound can live inside this region, so these vertical lines are useful reference values.
@@
In the figure, we have also indicated the values $\bar{a}=\pm 7.8(1000\,\mbox{km})^{2}$ as the vertical dashed green lines. Assuming $a+a^{\prime}$ is positive, then we should remain within the range by the WGC $|a|\lesssim 7.8(1000\,\mbox{km})^{2}$ or one of the polarizations would be inconsistent. (If $a+a^{\prime}$ is negative, we have no direct bound from WGC).
@@ VII Discussion
We have found an observational bound on the $R_{\mu\nu\alpha\beta}F^{\mu\nu}F^{\alpha\beta}$ coefficient of $|a|<26\,(1000\,\mbox{km})^{2}$. An important point is that the light propagation is polarization dependent. Ideally one would separate out the two polarizations and analyze each individually. However the data is not currently presented in this form. Instead we have exploited the fact that each individual data point has not shown a splitting into pairs (one raised and one lowered), or an enhanced variance, as would be expected if the higher dimension operator is altering the propagation, and hence we can still infer a bound on $a$.
Moreover, consider again the 2 eigenstates for this system: $\varepsilon=+1$ for polarization perpendicular to the plane of Cassini-earth sun and $\varepsilon=-1$ for polarization parallel to the plane of Cassini-earth-sun. Suppose we have one of these, say $\varepsilon=-1$ for polarization parallel to the plane to Cassini-earth sun, we anticipate that after reflecting off a mirror it will remain parallel. i.e., if the electric field is oscillating in such a plane, we see no reason it should rotate after hitting a mirror and returning. It can pick up a phase shift, but that does not alter the corrected time delay we are computing here. It will be reinforced in the round trip for a given polarization. However, if there is a mixture of polarizations, it is more subtle and possible cancellations can occur.
Our statistical analysis is overall a first pass, given that we do not have explicit polarization data. It would be useful in future work to search explicitly for a possible deviation in the behavior of each polarization. This may lead to a reduction in precision, or alternatively, a significantly improved bound (or a possible detection).